OpenStudy (anonymous):

How do you find an equation of the line that passes through the points (2,1) and (2,5)?

OpenStudy (anonymous):

I know the formulas; however I just can't seem to get the right answer...

OpenStudy (niel):

draw that line... in this case the answer will be clear as soon as you draw it

OpenStudy (anonymous):

I found the slope first which is4/0 = 0

OpenStudy (anonymous):

Then used y-y1=m(x-x1)...........No matter what, no correct answer..

OpenStudy (niel):

ok ... that is undefined.... there is no equation

OpenStudy (niel):

any fraction with a 0 on the bottom is impossible

OpenStudy (anonymous):

Okay, I put my answer as Undefined; however the teacher said it was wrong because I did not write an equation.

OpenStudy (anonymous):

I guess I need help plugging in the correct numbers into the Point Slope Form I wrote above. I keep getting it wrong.

OpenStudy (niel):

oh they want you to show your work ya gotcha.... y-y1=m(x-x1) (2,1) and (2,5) y-1=0(x-2)

OpenStudy (niel):

ask for a regrade on that problem

OpenStudy (anonymous):

Just curious, but I notice that in most of these examples, the y-1 always has a 1 ;however I thought we were plugging in a number from one of the points? Or is it always a 1, and then turns to just a y?

OpenStudy (niel):

y1 is just a label//// it meant the first y value same with x1

OpenStudy (anonymous):

Niel, thanks for the help

OpenStudy (anonymous):

the answer is x=2 the slope is undefined because its a vertical line

OpenStudy (anonymous):

I do not see how you get 2? With the formula, I get 4/0

OpenStudy (anonymous):

Which would equal, also undefined. How do you get 2? I am seriously confused with this..

myininaya (myininaya):

you have points (2,1) and (2,5) no matter what y is, x is always 2 (2,y) x=2

OpenStudy (anonymous):

I got...m=(5-1) / (2-2) = 4/0 = 0

OpenStudy (anonymous):

4/0 is undefined which indicates a vertical line plot both points on the graphand draw a line through them and it will be obvious

myininaya (myininaya):

slope is rise/run you have no run only rise

OpenStudy (anonymous):

a slope can be undefined in the case of a vertical line its just not a function using rectangular coordinates

myininaya (myininaya):

if you have a line going through (4,1) and (4,-2) then what would you say about the line (what is the equation of this line), JJSONFIRE?

OpenStudy (anonymous):

oh and \[4/0\neq0\]

OpenStudy (anonymous):

0 undefined..-3/0

myininaya (myininaya):

the slope undefined so again that means you have a veritcal (|), horizontal(-), or postive(/), or negative slope(\)?

OpenStudy (anonymous):

Okay, I understand that well now; however I am getting a little confused with writing the equations to indicate the points despite the undefined status. Teacher wants equations.

myininaya (myininaya):

that is a multiple choice question above that way? and the answer is vertical

OpenStudy (anonymous):

Myiniaya...Vertical

myininaya (myininaya):

right, both of are x's are the same (4,1) and (4,-2) so no matter what y is x will always be 4 so the equation is x=4

OpenStudy (anonymous):

Yes, vertical on a graph indicates it's undefined.

OpenStudy (anonymous):

horizontal line y= n slope is 0 vertical line x= n slope is undefined all others slope has a value that is an element of the reals as long as the \[\sqrt{-1}\] isn't involved

myininaya (myininaya):

so a line going through (2,1) and (2,-5), the equation is x=2 a line going through (-5,2) and (-5,5), the equation is x=-5 a line going through (-9,0) and (-9,3), the equation is x=-9 a line going through (2,3) and (4,3), the equation is y=3 a line going through (-4,5) and (-2,5) the equation is y=5 a line going through (-1,0) and (2,0) the equation is y=0

myininaya (myininaya):

in general, a line going through (a,b) and (a,c), the equation is x=a a line going through (g,h) and (f,h), the equation is y=h

myininaya (myininaya):

all other forms use point-slope formula or slope-intercept form

myininaya (myininaya):

y=h indicates a horiztonal line (slope=0) x=a indicates a veritcal line (slope=undefined) do you understand better?

OpenStudy (anonymous):

Yes, \much better..thanks everyone

OpenStudy (anonymous):

myiniaya, you've been a great big help..thanks.

myininaya (myininaya):

:)